OpenStudy (anonymous):

Which of the following is a solution to sqrtx-2 + 5 = 1? x= 6 x= 18 no solution x=16 which of the following is a solution to 4 sqrtx+3 = 12 x=0 x=3 x=6 x=12

OpenStudy (lalaly):

sqrt x-2=-4 square both sides x-2 = 16 x=18

OpenStudy (anonymous):

Second one: Divide both sides by 4. sqrt (x +3) = 3 Square both sides (whenever you square a square root, all you're left with is what was on the inside of the square root) x+3 = 3^2 x+3 = 9 x = 6